# Finding Maximum Independent Sets in Dynamic Graphs using Unsupervised Learning

### v1

@@ 5.1 Experimental Methodology
Ensuring that our experiments are as comprehensive as possible, we include both synthetically created as well as publicly available real-world datasets. Note that we build our dynamic graphs out of static graphs having similar characteristics. Specifically, we utilize their characteristic degree distribution as a basis to generate instances of graphs, which are then chained via edge events (additions or deletions) to form a dynamic graph. The synthetic graphs include Erdős–Rényi and Power Law graphs, while the real-world graphs are built using the TWITTER (Leskovec and Krevl, 2014), GERMANY (Orlowski et al., 2010), and BRAIN (Crossley et al., 2013) datasets. We select these real-world graphs due to their diverse topologies: TWITTER is a social network with dense small-world topology, BRAIN is a biological network (characteristic of sparse small-world topology), and GERMANY is a geological network with a mesh topology. The dynamic graphs we look at span a range of sizes—small (100 nodes, 50,000 time steps), medium (1,000 nodes, 100,000 time steps), and large (10,000 nodes, 5,000 time steps). Each dynamic graph is divided into its training, evaluation, and testing splits, arranged chronologically with increasing time steps. We use a 70:15:15 training, evaluation, testing split for small datasets and a 50:25:25 split for medium datasets. The large datasets only have 5,000 time steps because we use them solely for generalization experiments and do not train our model on them. Further details on dataset generation and dynamics will be added to the Appendix.
@@
We ran our complete pre-training and training pipeline using the two model variants: BCAS and No-CAS. For our experiments with the BCAS variant, we set $\alpha=\beta=0.25\cdot\text{diam}(G_{0})$, i.e., half of the radius of the first snapshot of the dynamic graph.
To keep our comparisons balanced and complete, we compare against a variety of existing works for finding Maximum Independent Sets. While we compare to the work by Zheng et al. (2019), which is designed for dynamic graphs, all the other works we consider are intended for static graphs. For these, we compute the solution for each snapshot in the dynamic graph independently.
  • One of the rule-based update algorithms for maintaining MaxIS in dynamic graphs, referred to as Update-Algo moving forward (Zheng et al., 2019).
  • Current state-of-the-art GNN-based machine learning model for MaxIS in static graphs: comparator model by Brusca et al. (2024), referred to as DP-GNN in this work.
  • As an approximate heuristic search method, we run the deterministic greedy algorithm, referred to as Greedy-MaxIS in the results.
  • Gurobi, the most commonly used and commercially available mixed integer programming solver (Gurobi Optimization, LLC, 2024).
The reported results for our model variants correspond to the best performance across three runs with different random seeds. We train and test our model variants and all the other algorithms on a AMD EPYC 7413 CPU with single A100:20gb GPU machine. For the learning methods, training was halted after seven days; for Gurobi, we enforced a time limit of 1 s.
@@
We compare our model against the different methods on the specified datasets using three metrics: approximation ratios, peak memory usage, and per-graph solution generation time. The approximation ratio (referred to as performance) refers to the factor by which an algorithm’s solution deviates from the optimal solution of a problem. In our case, it is defined as the ratio of the MaxIS solution obtained by a method to the corresponding solution produced by Gurobi, which we use as our baseline. Peak memory usage denotes the maximum memory consumed by each method while training on a given dataset. Per-graph solution generation time (referred to as runtime) refers to the time taken by a method to process a single snapshot and generate a solution. We report the average approximation ratio and runtime (in seconds per graph, s/g) over the entire test dataset, which consists of a sequence of snapshots (time steps) for each dynamic graph.
@@ 5.2 Results
The results draw out the efficient scalability and performance of our model, where both our model variants outperform DP-GNN across all the medium-sized datasets (Table 2), both in terms of performance (1.65.4x better) and runtime (2001,220x faster). DP-GNN gets handstrung by its autoregressive approach, and its enormously slow runtime even prevented us from evaluating it on the entire test dataset. Adding on to that, we are much more memory efficient and consume 11.527.6x less memory in comparison. Even when compared on small-sized datasets (Table 1), our model variants perform fairly well while the BCAS variant is 1.32.6x faster in runtime.
Our BCAS variant compares favourably to Greedy-MaxIS in terms of performance on the BRAIN dataset, while having almost the same runtime. While only slightly worse in performance on the ER and PL datasets, we are still 24.8x faster in runtime. We draw out this observation to argue that our method can perform better than a Greedy-MaxIS method, and Greedy-MaxIS is not necessarily a better option as posited by Angelini and Ricci-Tersenghi (2023).
On the medium-sized datasets, we almost always outperform Gurobi in at least one of our variants. At the same time, our runtime is only a fraction of a second compared to Gurobi’s 1s time limit.
@@
We highlight our model’s generalization ability by using our model variants trained on small-sized synthetic datasets to evaluate on the corresponding large-sized datasets (Table 3), which are 100x largerthan small-sized datasets. We observe that the No-CAS variant scales fairly in runtime while achieving competitive performance relative to the Update-Algo (0.89x accurate on ER and matching the performance in PL). It is able to surpass Greedy-MIS in terms of runtime (1.55x faster) while being competitive in performance (0.930.99x accurate). Curiously, BCAS is 23x faster than Greedy-MaxIS in the PL dataset with nearly matched performance. We compared against Gurobi capped at 10 seconds and see that our model exceeds its performance while being faster on the ER dataset and matching the performance on the PL dataset. Once again, the stark contrast between DP-GNN and a single-shot method like ours is evident; there is a 4 orders of magnitude difference in solution generation time per graph, and it is prohibitively long to complete even a portion of the dataset and report conclusive performance values.
Table 1: Test dataset performance (w.r.t. Gurobi) for small-sized datasets. We report the means and the std. deviations of the performance along with the runtime (in seconds per graph), averaged across the test dataset. We report the peak memory utilized while training by learning methods in GigaBytes (GB). We separate the learning based methods from rule-based methods and highlight the learning methods with the best performance (bold-black) and the best runtime (blue) for each dataset.
Method ($\downarrow$) Dataset ($\rightarrow$) ER PL TWITTER GERMANY
@@
$0.83\pm 0.050$
$(0.030s/g)$
($2.3$ GB)
$0.98\pm 0.015$
$(0.035s/g)$
($3$ GB)
$0.94\pm 0.030$
$(0.063s/g)$
($3$ GB)
$0.92\pm 0.046$
$(0.046s/g)$
($2$ GB)
No-CAS
$0.93\pm 0.030$
$(0.075s/g)$
($3.5$ GB)
$0.88\pm 0.050$
$(0.068s/g)$
($3.4$ GB)
$0.94\pm 0.030$
$(0.087s/g)$
($2.6$ GB)
$0.94\pm 0.042$
$(0.061s/g)$
($2$ GB)
DP-GNN
$\mathbf{0.98\pm 0.058}$
$(0.079s/g)$
($3.7$ GB)
$\mathbf{0.99\pm 0.051}$
$(0.086s/g)$
($3.6$ GB)
$\mathbf{0.97\pm 0.036}$
$(0.080s/g)$
($6.4$ GB)
$\mathbf{0.99\pm 0.035}$
$(0.071s/g)$
($1.9$ GB)
Update-Algo
$0.99\pm 0.012$
$(0.004s/g)$
$0.99\pm 0.012$
$(0.006s/g)$
$0.99\pm 0.009$
$(0.005s/g)$
$0.99\pm 0.017$
$(0.005s/g)$
Greedy-MaxIS
$0.98\pm 0.058$
$(0.002s/g)$
$0.99\pm 0.051$
$(0.002s/g)$
$0.97\pm 0.039$
$(0.002s/g)$
$0.98\pm 0.073$
$(0.001s/g)$
Table 2: Test dataset approximation ratios (w.r.t. Gurobi) ran for medium-sized datasets. We report the results in similar manner as Table 1
Method ($\downarrow$) Dataset ($\rightarrow$) ER PL BRAIN
BCAS
$1.00\pm 0.025$
$(0.023s/g)$
($2.1$ GB)
$0.96\pm 0.010$
$(0.076s/g)$
($2.2$ GB)
$\mathbf{1.41\pm 0.055}$
$(0.054s/g)$
($6.1$ GB)
No-CAS
$\mathbf{1.08\pm 0.025}$
$(0.270s/g)$
($3$ GB)
$\mathbf{0.99\pm 0.005}$
$(0.250s/g)$
($2.3$ GB)
$1.34\pm 0.049$
$(0.498s/g)$
($4.7$ GB)
DP-GNN
$0.45\pm 0.02$
$(28.08s/g)$
($44$ GB)
$0.59\pm 0.022$
$(33.45s/g)$
($26.4$ GB)
$0.26\pm 0.059$
$(10.97s/g)$
($129.7$ GB)
Update-Algo
$1.19\pm 0.022$
$(0.006s/g)$
$1.00\pm 0.001$
$(0.006s/g)$
$1.49\pm 0.055$
$(0.014s/g)$
Greedy-MaxIS
$1.15\pm 0.022$
$(0.111s/g)$
$1.00\pm 0.000$
$(0.165s/g)$
$1.40\pm 0.050$
$(0.045s/g)$
Table 3: Generalization results for large-sized datasets ($10000$ nodes), reported similarly as Table 1. Note, we do not specific the peak memory usage as no bespoke training was executed.
Method ($\downarrow$) Dataset ($\rightarrow$) ER PL
BCAS
$0.63\pm 0.085$
@@
$(36239.508s/g)$
$(39994.408s/g)$
Update-Algo
$1.31\pm 0.027$
$(0.029s/g)$
$0.99\pm 7.611$
$(0.02s/g)$
Greedy-MaxIS
$1.25\pm 0.025$
$(8.373s/g)$
$1.00\pm 0.000$
$(16.153s/g)$
Gurobi(10s)
$1.04\pm 0.026$
$(10.000s/g)$
$1.00\pm 0.000$
$(0.050s/g)$
@@ 6 Conclusion
We present a single-shot, unsupervised learning model based on Graph Neural Networks (GNNs) for solving the Maximum Independent Set (MaxIS) problem in dynamic graphs, where the topology evolves over time. Unlike heuristic learning approaches, our method learns an update mechanism analogous to those used in traditional rule-based algorithms for dynamic combinatorial optimization. We conduct experiments across various graph types and edge dynamics, demonstrating that our model outperforms the DP-GNN model (Brusca et al., 2024) for static graph MaxIS in terms of scalability in both runtime and memory, while maintaining strong approximation performance even on dynamic graphs with up to 10,000 nodes. In the future, we plan to extend our model into a general framework for learning update algorithms applicable to a broader class of combinatorial optimization problems.
@@
While our experiments suggest that our model may eventually outperform Greedy-MaxIS in terms of both runtime and accuracy, we need to conduct extensive evaluations to conclusively demonstrate this. Similarly, we only run experiments for small–medium real-world graphs; it would be interesting to see how our model performs on large real-world dynamic graphs. Finally, the model currently makes limited use of the temporal dynamics of the dynamic graphs to learn an update mechanism, specifically looking at the effect of a single edge event. Future work is needed to leverage longer temporal dependencies and improve the update mechanism further.

### v2

@@ 5.1 Experimental Methodology
We evaluate model performance across both synthetic and empirical dynamic graphs. For synthetic graphs, we consider Erdős–Rényi (ER) and Power Law (PL) graphs with randomly generated initial structures. For empirical graphs, we used initial structures given by GERMANY (Orlowski et al., 2010), a geographical network with a mesh topology; TWITTER (Leskovec and Krevl, 2014), a social network subgraph with a dense small-world topology; and BRAIN (Crossley et al., 2013), a biological network with a sparse small-world topology. Across all graphs, dynamics are produced by sampling edge additions or deletions according to the initial snapshot’s degree distribution for a desired number of time steps.
The dynamic graphs we consider span a range of sizes: small ($\approx$100 nodes, 50,000 time steps), medium ($\approx$1,000 nodes, 100,000 time steps), and large ($\approx$10,000 nodes, 5,000 time steps). Each dynamic graph is divided into its training, evaluation, and testing splits, arranged chronologically with increasing time steps. We use a 70:15:15 training, evaluation, testing split for small datasets and a 50:25:25 split for medium datasets. The large datasets only have 5,000 time steps because we use them solely for generalization experiments and do not train our model on them.
@@
We ran our complete training pipeline using the two model variants: BCAS and No-CAS. For the BCAS variant, we set $\alpha=\beta=0.25\cdot\text{diam}(G_{0})$, i.e., half of the radius of the first snapshot of the dynamic graph. For each variant, we performed three training runs with different random seeds and report results for the median run. Since ours is the only learning method for MaxIS in the dynamic setting, we compare our method against a commercial mixed integer programming solver and existing learning methods for MaxIS on static graphs. Since these methods are intended for static graphs, we apply them to each snapshot of a dynamic graph independently.
  • Gurobi (Gurobi Optimization, LLC, 2024) is the most commonly used and commercially available mixed integer programming solver. We use Gurobi’s default settings for mixed integer program solving, which according to their documentation, is a branch-and-bound search.
  • DP-GNN (Brusca et al., 2023) is the state-of-the-art learning model for MaxIS in static graphs, combining dynamic programming with GNNs to form a multi-step, autoregressive comparator model.
  • Erdős-GNN (Karalias and Loukas, 2020) is an earlier GNN-based, single-shot, probabilistic learning model for combinatorial optimization in static graphs.
  • LwDMIS (Ahn et al., 2020) is a deep reinforcement learning method that adaptively adjusts its stages and defers decisions about nodes’ MaxIS membership to improve scalability.
All methods were trained (if applicable) and tested on a machine with an AMD EPYC 7413 CPU and a 20 GB slice of one NVIDIA A100 GPU. For the learning methods, training was halted after seven days. For Gurobi, we enforced a time limit of 1 s per snapshot; this is sufficient for producing exact solutions for small datasets, but yields approximations for medium and large ones.
@@
We evaluate each method using three metrics: approximation ratio (performance), per-graph solution generation time (runtime), and peak memory usage. Performance refers to the mean ratio of a method’s candidate MaxIS size against that of Gurobi’s across all snapshots in a dynamic graph. Runtime refers to the mean time taken by a method to process a single snapshot and generate a candidate MaxIS across all snapshots in a dynamic graph (reported as seconds per graph, s/g). Peak memory usage denotes the maximum memory used by a method during training.
@@ 5.2 Results
(a) Performance
(b) Runtime
(c) Peak Memory Usage
Figure 2: Comparisons of models’ (a) performance, (b) runtime, and (c) peak memory usage across the seven small and medium datasets. The best method for each metric and dataset is marked with a ($\ast$); for performance, all methods for a given dataset whose mean performances are within one standard deviation (error bars) of the highest mean performance are considered “best”. Notably, DP-GNN is very slow on larger graphs, so its results for the medium datasets only reflect evaluation on the first 5,000 of 25,000 total testing timesteps. Also, we observed LwDMIS fail to train on some graph snapshots as its search space collapsed to null; thus, we only report peak training memory usage for those methods that successfully completed the entire training process.
The results highlight the efficient scalability and overall performance of our model. Both the BCAS and No-CAS variants of our model consistently outperform other learning-based methods in terms of runtime and peak memory usage, while also achieving comparable or superior performance across all graph sizes. Surprisingly, DP-GNN—although the current state-of-the-art—performs well only on small datasets. On the medium datasets, its performance is 1.30–4.20x worse than any next best method across datasets (Figure 2(a)) while running more than 1,000x slower (Figure 2(b)) and consuming up to 27.44x more memory (Figure 2(c)) than other methods. In contrast, earlier methods such as LwDMIS and Erdős-GNN exhibit better tradeoffs over different graph sizes: LwDMIS is reasonably fast but performs similarly or worse than Erdős-GNN, especially on small graphs, while Erdős-GNN achieves its stronger performance at the cost of slower speeds and worse scalability. Our model variants strike an effective balance between these, delivering 1.071.22x better performance than either method (Figure 2(a)) while also reducing runtime (6.1824.8x faster, Figure 2(b)) and saving up to 2.7812.97x memory (Figure 2(c)). This demonstrates the efficiency and scalability of our model when compared to any other existing learning method.
BCAS vs. No-CAS
Table 1: Pearson correlation for model performance and runtime vs. properties of graph snapshots. Correlation coefficients are calculated over all 75,000 graph snapshots in the testing splits of our medium datasets. Due to the large number of samples, we set our threshold for significance at $p<10^{-10}$. All reported correlations are significant with $p<10^{-40}$, except the one value marked with a ($\dagger$), which has $p=\text{0.00095}$.
Performance Runtime
Graph Properties BCAS No-CAS BCAS No-CAS
# Edges 0.99 0.97 –0.09 0.98
Min. Degree 0.91 0.94 –0.15 0.92
Max. Degree 0.96 0.89 0.12 0.92
Ave. Degree 0.99 0.97 –0.06 0.98
Ave. Clustering Coefficient 0.99 0.95 –0.01 0.97
Diameter –0.68 –0.79 0.43 –0.74
Ave. Node Distance –0.75 –0.85 0.39 –0.81
Table 1 shows correlations between our BCAS and No-CAS models’ performance and runtime vs. properties of the graph snapshots in the testing datasets. For performance, the same graph properties help and hurt both BCAS and No-CAS: both models perform better on graphs with many edges, high-degree nodes, and dense clustering but perform worse when diameter and node distance increase. This is intuitive, as both variants aggregate information from only their local neighbors and thus benefit from dense structure but suffer from distant changes that are not propagated. For runtime, there is a sharp contrast between the variants. BCAS runtimes are largely uncorrelated with graph properties—owing to its strictly local cascades and aggregation—exhibiting only a weak positive correlation with diameter and distance (stemming from the fact that $\alpha=\beta$ are set as a fixed fraction of diameter). No-CAS runtimes, on the other hand, are strongly determined by density (i.e., edge counts, degrees, and clustering), since every node in the graph must perform local aggregation at every time step and the complexity of local aggregation is entirely dependent on neighborhood size.
@@
We highlight our model’s generalization ability by training our model variants on small synthetic datasets and then evaluating them on corresponding datasets 100x their size (Table 2). DP-GNN, Erdős-GNN, and LwDMIS display the same strengths and weaknesses as for the smaller datasets: DP-GNN scales so poorly that its runtime is four orders of magnitude worse than the others, Erdős-GNN achieves decent performance but is slower than the remaining methods, and LwDMIS runs very quickly at the cost of inferior performance. Our No-CAS variant, although not as fast as LwDMIS, achieves far superior performance to any other learning method and even ties Gurobi’s performance in a third of the runtime. On the PL dataset specifically, BCAS achieves the same great performance, but is 4.97x faster than No-CAS and 14.73x faster than Gurobi.
Table 2: Generalization results for large-sized datasets (10,000 nodes, 5,000 time steps). Performance is reported as the mean and standard deviation of MaxIS sizes across snapshots against those of Gurobi’s (run for 10 s). Runtime is reported in parentheses as mean seconds per graph snapshot. Notably, DP-GNN is prohibitively slow on these large graph snapshots, taking nearly a week just to evaluate the first six snapshots of each dataset. Thus, we report performance only for those methods that could be evaluated on the entire datasets.
Dataset ($\rightarrow$)
Method ($\downarrow$)
ER
($n=$ 10,000)
PL
($n=$ 10,000)
@@
0.63 $\pm$ 0.085
@@
N/A
(36,239.508 s/g)
N/A
(39,994.408 s/g)
Erdős-GNN
1.07 $\pm$ 0.023
(19.202 s/g)
0.94 $\pm$ 0.009
(18.415 s/g)
LwDMIS
0.99 $\pm$ 0.020
(0.541 s/g)
0.92 $\pm$ 0.003
(0.296 s/g)
@@ 6 Conclusion
We present a single-shot, unsupervised learning model based on graph neural networks (GNNs) for the Maximum-Independent-Set (MaxIS) problem in dynamic graphs, where the topology evolves over time. Unlike heuristic learning approaches, our method learns an update mechanism analogous to those used in traditional rule-based algorithms for dynamic combinatorial optimization. We conduct experiments across various dynamic graph topologies, demonstrating that our model outperforms the DP-GNN (Brusca et al., 2023) and Erdős-GNN (Karalias and Loukas, 2020) models for MaxIS in terms of both runtime and memory scalability while maintaining strong approximation ratios even on dynamic graphs with up to 10,000 nodes. The prior LwDMIS model Ahn et al. (2020), although faster, lags far beyond our model’s performance. In the future, we plan to extend our model into a general framework for learning update algorithms applicable to a broader class of combinatorial optimization problems.
@@
Our experiments demonstrate that our model outperforms existing learning methods for MaxIS in terms of approximation ratios, runtime, and training memory. However, a full comparison to traditional, non-learning heuristic methods is beyond the scope of this paper. Some preliminary experiments suggest that on large dynamic graphs, our model is faster than and achieves comparable performance to the iterative application of a greedy method—assuaging criticisms of learning-based approaches such as (Angelini and Ricci-Tersenghi, 2022)—but deterministic rule-based update algorithms that explicitly consider the dynamic setting may perform even better (Zheng et al., 2019; Gao et al., 2022). Also, the dynamics for all our datasets—synthetic or empirical—are sampled synthetically (though in a degree distribution-preserving way). It would be interesting to evaluate the performance of these methods on large dynamic graphs whose structure and dynamics are both empirical. Finally, our model currently makes limited use of temporal dynamics when learning its update mechanism, looking only at the effect of a single edge event. Future work is needed to leverage longer temporal dependencies and improve the update mechanism further.
Appendix A Model Pseudocode
Algorithms 13 show pseudocode for our model’s initial MaxIS generation procedure, update training process, and integral solution generation procedure, respectively. Throughout, graph snapshots $G_{t}$, edge events $\mathcal{E}_{t}$, signals $s_{t}(\cdot)$, node memories $m_{t}(\cdot)$, node estimates $p_{t}(\cdot)$, and node losses $\ell_{t}(\cdot)$ are defined as they were in the main text. Bold versions of these notations refer to the vector of all node properties of that type (e.g., $\mathbf{m}_{t}$ refers to all node memories). In a slight abuse of notation, we use $\mathcal{N}_{t}^{x}(\mathcal{E}_{t})$ to denote all nodes within an $x$ hop-distance from the edge event $\mathcal{E}_{t}$; i.e., the endpoints of the edge event are in $\mathcal{N}_{t}^{0}(\mathcal{E}_{t})$, those endpoints and their neighbors are in $\mathcal{N}_{t}^{1}(\mathcal{E}_{t})$, and so on.
Algorithm 1 Initial MaxIS Generation Epoch
1:Initialize all node memories as $\mathbf{m}_{0}\leftarrow\mathbf{0}$.
2:for all edges $(u,v)\in E_{0}$ do
3:  Let $\mathcal{E}$ be the edge event of adding $(u,v)$.
4:  Do event handling with signal $s\leftarrow[\texttt{enc}(\mathcal{E})~||~0]$ as in Eq. 1.
5:  Update $m_{0}(u)\leftarrow\texttt{GRU}([m_{0}(u)~||~s])$ and $m_{0}(v)\leftarrow\texttt{GRU}([m_{0}(v)~||~s])$ as in Eq. 2.
6:Compute node estimates $\mathbf{p}_{0}$ from the updated node memories $\mathbf{m}_{0}$ as in Eqs. 35.
7:Compute node losses $\mathbf{\ell}_{0}$ from $\mathbf{p}_{0}$ as in Eq. 6.
8:Compute cumulative loss $L_{0}\leftarrow\sum_{v\in V}\ell_{0}(v)$.
Algorithm 2 Update Training Epoch
1:Initialize node memories $\mathbf{m}_{0}$ using Algorithm 1.
2:for all edge events $\mathcal{E}_{t}\in(\mathcal{E}_{1},\mathcal{E}_{2},\ldots,\mathcal{E}_{T})$ in the training set do
3:  for all nodes $v\in\mathcal{N}_{t}^{\alpha}(\mathcal{E}_{t})$ do
4:   Do event handling with $s_{t}(v)\leftarrow[\texttt{enc}(\mathcal{E}_{t})~||~r_{t}(v)]$ as in Eq. 1.
5:   Update $m_{t}(v)$ using $s_{t}(v)$ as in Eq. 2.
6:  for all nodes $v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})$ do
7:   Compute estimate $p_{t}(v)$ from $m_{t}(v)$ as in Eqs. 35.
8:  for all nodes $v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})$ do
9:   Compute node loss $\ell_{t}(v)$ from $\mathbf{p}_{t}$ as in Eq. 6.
10:  Compute cumulative loss $L_{t}\leftarrow\sum_{v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})}\ell_{t}(v)$.
Algorithm 3 Integral Solution Generation (Testing & Inference)
1:Initialize node memories $\mathbf{m}_{0}$ using Algorithm 1.
2:for all edge events $\mathcal{E}_{t}\in(\mathcal{E}_{1},\mathcal{E}_{2},\ldots,\mathcal{E}_{T})$ in the testing set do
3:  for all nodes $v\in\mathcal{N}_{t}^{\alpha}(\mathcal{E}_{t})$ do
4:   Do event handling with $s_{t}(v)\leftarrow[\texttt{enc}(\mathcal{E}_{t})~||~r_{t}(v)]$ as in Eq. 1.
5:   Update $m_{t}(v)$ using $s_{t}(v)$ as in Eq. 2.
6:  for all nodes $v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})$ do
7:   Compute estimate $p_{t}(v)$ from $m_{t}(v)$ as in Eqs. 35.
8:  for all nodes $v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})$ do $\triangleright$ Estimate Rounding
9:   if $p_{t}(v)\geq 0.5$ then $I_{t}(v)\leftarrow 1$.
10:   else $I_{t}(v)\leftarrow 0$.
11:  while there exists a node $v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})$ with violations do $\triangleright$ Violation Removal
12:   Choose the node $v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})$ with the most violations.
13:   Remove $v$ from the MaxIS wtih $I_{t}(v)\leftarrow 0$.